# The First and Second-Order Asymptotics of Covert Communication over AWGN Channels

### v4

@@ IV-C Existence of the Codebooks and Achievability Bound
Our result on the size of key bits is slightly different from the case in discrete memoryless channels [23], where the key bits is $O(1)$ if the adversary is less noisy ($D_{Q}\leq D_{P}$). In this setting, the number of key bits also depends on the parameter of truncation $\mu$ with finite blocklength over AWGN channels. Even if the channel between Alice and Willie is severer than the main channel, the number of key bits should be of $O(\sqrt{n})$ if the noise level is only a bit larger. This is the necessary cost of power constraint for the sending codewords with finite blocklength over AWGN channels.
IV-D An Upper Bound on the Achievable $M_{n}$ and A Lower Bound on $K_{n}$
Our previous analysis is based on the fact the distribution $\bar{P}_{n}^{n}$ and $P_{n}^{n}$ are close when $n$ is large and the parameter of truncation $\mu$ is fixed. Although it will lead to achievable throughput and the number of key bits, these bounds may be not tight because the choice of $\mu$ as a fixed constant may not be optimal in the asymptotic sense. In the following, we revisit the analytic formula of $\Delta$ in (23) and seek for proper $\mu$ by numerical computation. Especially, we consider $\mu$ be in the form of $\mu(n)=1-\frac{1}{n^{\eta}}$ where $\eta>0$ and re-investigate the first and second order asymptotics of our previous results.
The following Fig. 3 shows the numerical behavior of $\Delta$ with different values of $\eta$. It is obvious that $\eta\in(0,0.2]$ still leads to $\Delta\rightarrow 0$ as $n\rightarrow\infty$. In this case,
$\,\ \ \ \,\mu(n)\rightarrow 1$ (123)
and
$\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))\rightarrow 0,\,\ \ \ \,\hat{\mathsf{C}}(\Psi,\mu\Psi(n),\sigma^{2})\rightarrow 0$ (124)
as $n\rightarrow\infty$. We have an upper bound on the achievable $M_{n}$:
$\begin{split}\sup_{\mu}&\{\log M_{n}:M_{n}\,\,\text{is achievable}\,\}\\ \leq&nC(\Psi(n))-\sqrt{n\mathsf{V}(\Psi(n),\Psi(n))}\\ \times&\Phi^{-1}(1-\epsilon_{1})+O(\log n),\end{split}$ (125)
where $\Psi(n)$ satisfies
@@ IV-D An Upper Bound on the Achievable $M_{n}$ and A Lower Bound on $K_{n}$
$+$ $\mathbb{D}(\bar{P}_{n}^{n}\|P^{n}_{n})+(1+\iota^{\prime})\mathbb{D}(\bar{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{0})+\lambda_{3}\leq\delta\}$
[…]
$=$ $\Psi(\delta,n)$ (126)
where $\Psi(\delta,n)$ is the solution of
$\mathbb{D}(\bar{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{0})=\delta\,\ \ \,\text{with}\,\ \,\bar{Q}^{(n)}_{1}\sim\mathcal{N}(\bm{0},(\sigma^{2}+\Psi(n))\bm{I}_{n}).$ (127)
Therefore, we arrive at an upper bound of $\log M_{n}$
$\begin{split}\sup_{\mu}&\{\log M_{n}:M_{n}\,\,\text{is achievable}\,\}\\ \leq&\frac{n}{2}\log\big{(}1+\Psi(\delta,n)\big{)}-\sqrt{n\mathsf{V}(\Psi(\delta,n))}\\ \times&\Phi^{-1}(1-\epsilon_{1})+O(\log n)\end{split}$ (128)
where $\mathsf{V}(P)=\frac{P(2P+1)}{2(1+P)^{2}}\log^{2}e$.
From (117), we have a lower bound of the number of bits $K_{n}$ as
$\begin{split}\inf_{\mu}\log K_{n}\geq\begin{cases}\frac{n}{2}\log\big{(}1+\frac{\Psi(\delta,n)(1-\sigma^{2})}{\sigma^{2}(1+\Psi(\delta,n))}\big{)}+\sqrt{n\mathsf{V}(\Psi(\delta,n))}\\ \,\ \ \ \ \,\times\Phi^{-1}(1-\epsilon_{1})\,\ \ \ \ \ \ \ \ \ \ \ \,\text{if}\,\ \,\sigma^{2}<1,\\ 0\,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,otherwise.\\ \end{cases}\end{split}$ (129)
Note that the above bounds represent the supreme of the achievable throughputs and the infimum of the key bits by neglecting the effect of truncation in the random coding scheme. However, in practice, truncation is necessary in the random coding with finite blocklength because we can not perfectly simulate Gaussian distribution by signals, which means the bounds can not be achieved by our coding scheme with finite $n$. Nevertheless, the transmitter can choose proper $\mu(n)$ in Theorem 3 to approach them, and we believe that when $n$ is sufficiently large, the gap between the achievable throughput in (98) and (128) is small. Besides, the above bound also serves as a comparison with the converse bound to show the optimality of the random coding in the achievability, which will be further illustrated later.
Fig. 3: Numerical behavior of $\Delta$ with proper chosen $\eta$.
V Converse Bound under Covert Constraint
In this section, the converse bound under covert constraint (9) is investigated, and the technique is to reconcile the bound in terms of the output distribution of the channel to get an upper bound of average power level of the codebook. The intuitive idea stems from the practice of communication engineering where the engineers know to blend the signals with the noise by using signals with sufficiently low power. Our technique is similar as [23] in binary-input discrete memoryless channel, where upper bounds of the weight of the codewords are obtained to get the upper bounds of the throughput.
For the definition of $M^{*}(n,\epsilon,\delta)$, we are considering an upper bound of the maximal throughput of all coding schemes satisfying covert constraint (9). The tightest upper bound corresponds the solution of the following optimization problem:
$\begin{split}\underset{P_{X^{n}}}{\max}&\,\ \,M(P_{X^{n}})\\ \,\ \ \,s.t.\,\ \ \,&\mathbb{D}(\hat{Q}^{(n)}_{\bm{Z}},\hat{Q}^{(n)}_{0})\leq\delta.,\end{split}$ (130)
where $P_{X^{n}}=\frac{1}{KM}\sum_{s=1}^{K}\sum_{w=1}^{M}\bm{x}_{sw}$ is the distribution of the code in $\mathbb{R}^{n}$, $M(P_{X^{n}})$ is the throughput of the code. Nevertheless, the solution of the above optimization problem for each $n$ is generally unknown. Instead, we relax the condition on the the form of the distribution $P_{X^{n}}$, and consider all distributions with maximal second order moment constraint.
First, we define the distribution of any code as $P_{\bm{x}}$ to incorporate all random coding schemes (including deterministic coding as a special case), and regard each codeword as a vector generated by $P_{\bm{x}}$. The average second moment of the distribution is expressed as $\frac{1}{n}Tr(Cov(\bm{x}))$, where $Cov(\bm{x})$ is the covariance matrix of the random codeword $\bm{x}$ and $Tr$ denotes its trace. For example, when we consider Gaussian random codes with sufficiently large length $n$ and average power $n\mu\Psi(n)$, we have $P_{\bm{x}}\sim\mathcal{N}(\bm{0},\mu\Psi(n)\bm{I}_{n})$ and the covariance matrix is $\mu\Psi(n)\bm{I}_{n}$.
Below we introduce some distributions of the codes for covert communication over the AWGN channels.
  • Let $\mathbb{P}^{(n)}_{in}$ be any distribution of $\bm{x}$ under given second moment constraint
    $\frac{1}{n}Tr(Cov(\bm{x}))=\Psi_{a}$ (131)
    and the induced output distribution of the AWGN channel (with background noise $\mathcal{N}(0,\sigma^{2})$) is denoted as $\mathbb{Q}^{(n)}_{in}$.
  • Let $\mathbb{P}^{(n)}_{G}$ be the $n$ dimensional Gaussian distribution with i.i.d coordinates, and each coordinate has zero mean and variance $\Psi_{a}$. The induced output distribution of the AWGN channel (with background noise $\mathcal{N}(0,\sigma^{2})$) is $\mathbb{Q}^{(n)}_{G}\sim\mathcal{N}(\bm{0},(\sigma^{2}+\Psi_{a})\bm{I}_{n})$.
  • Let ^$\hat{\mathbb{P}}^{(n)}_{G}$ be any $n$ dimensional Gaussian distribution with zero mean and covariance matrix $\bm{K}$ which satisfies $\frac{1}{n}Tr(\bm{K})=\Psi_{a}$. The induced output distribution of the AWGN channel is ^$\hat{\mathbb{Q}}^{(n)}_{G}$ with covariance matrix $\bm{K}+\sigma^{2}\bm{I}_{n}$.
For these distributions with the same parameter $\Psi_{a}$, we consider the particular discrimination - KL divergence between these distributions and $\hat{Q}^{(n)}_{0}$. For the case of $n=1$, it is shown in [14] (Theorem 5) and [9] (Theorem 2) that with zero-mean Gaussian distribution $\mathbb{P}^{(1)}_{G}=\mathcal{N}(0,\Psi_{a})$ is the minimizer of $D(\mathbb{Q}^{(1)}_{in}\|\mathbb{Q}^{(n)}_{0})$. For $n\geq 1$, we have the following generalization, the proof of which can be found in Appendix E.
Theorem 4.
For the above distributions $\mathbb{P}^{(n)}_{in}$, $\mathbb{P}^{(n)}_{G}$ and ^$\hat{\mathbb{P}}^{(n)}_{G}$, we have
$\forall n,\,\ \,D(\mathbb{Q}^{(n)}_{in}\|\mathbb{Q}_{0}^{(n)})\geq D(\hat{\mathbb{Q}}^{(n)}_{G}\|\mathbb{Q}_{0}^{(n)})\geq D(\mathbb{Q}^{(n)}_{G}\|\mathbb{Q}_{0}^{(n)}).$
Theorem 4 implies the following fact. For any randomly generated code with fixed second moment condition (131), Gaussian random codes minimize KL divergence between the induced output distribution and the distribution of the noise at the adversary. The equivalent statement is as follows.
Corollary 2.
Any codebook with distribution $\mathbb{P}^{(n)}_{in}$ which satisfies the covert constraint, then the corresponding $\Psi_{a}$ in (131) should be less than the solution $x$ of the equation
$D(\bar{\mathbb{Q}}^{(n)}_{G}\|\mathbb{Q}^{(n)}_{0})=\delta,$ (132)
where ¯$\bar{\mathbb{Q}}^{(n)}_{G}=\mathcal{N}(\bm{0},(\sigma^{2}+\varpi)\bm{I}_{n}).$
Theorem 5.
In AWGN channels with channel models in (5) and (6), Any codebook under covert constraint (9) should satisfy
$\log M^{*}(n,\epsilon,\delta)$ $\leq\frac{n}{2}\log\big{(}1+\Psi(\delta,n)\big{)}-\sqrt{n\mathsf{V}(\Psi(\delta,n))}$
$\times$ $\Phi^{-1}(1-\epsilon_{1})+O(\log n)$ (133)
Proof.
For any codebook, the second moment is equivalent to the average power since it is assumed that the codewords are equiprobable. By solving the equation (132), we get the variance $\Psi^{\star}(n)$ of $\mathbb{P}^{(n)}_{G}$. From the arguments around (127), $\Psi^{\star}(n)=\Psi(\delta,n)$. Hence, for any other distribution with variance $\Psi_{a}$, we have $\Psi_{a}\leq\Psi(\delta,n)$ to satisfy the covert constraint $D(\mathbb{Q}^{(n)}_{in}\|\mathbb{P}_{0}^{(n)})\leq\delta$. In other words, Gaussian random codes will have largest average power among all random codes under covert constraint (9), i.e., we have
$M^{*}(n,\epsilon,\delta)\leq M^{*}_{a}(n,\epsilon,\Psi(\delta,n)).$ (134)
Consequently,
$\log M^{*}(n,\epsilon,\delta)$
$\leq\log M^{*}_{a}(n,\epsilon,\Psi(\delta,n))$
$\overset{(a)}{\leq}\log\left[\frac{1}{1-\frac{n}{n+1}}M^{*}_{m}\left(n,\epsilon,(1+\frac{1}{n})\Psi(\delta,n)\right)\right]$ (135)
In step (a), we have used
$M^{*}_{a}(n,\epsilon,P)\leq\frac{1}{1-\frac{P}{P^{\prime}}}M^{*}_{m}(n,\epsilon,P^{\prime}),$ (136)
from Lemma 39 in [2] by letting $P^{\prime}=\frac{n+1}{n}P$. After some algebra, we finally get (5). ∎
@@ V Converse Bound under Covert Constraint
Note that the converse bound is the same as (128). It shows that the coding scheme in the achievability is optimal. In practice, the maximal throughput can be obtained by choosing proper parameter $\mu$, which depends on the blocklength $n$ and the noise level $\sigma^{2}$ at the adversary.
@@
VI A Local Information Geometry Viewpoint of Gaussian Random Coding for Covert Communication
In Section IV, truncated Gaussian distribution is adopted as the generating distribution of the AWGN covert codes. From the arguments in Section IV-D and Section V, the distribution is optimal in the sense that it leads to nearly maximal coding rate when $n$ is large. In this section, we analyze it from a viewpoint of local information geometry, which helps us to understand the underlying principle of choosing generating distribution for random coding in covert communication.
VI-A Quasi-$\varepsilon$-neighborhood Distributions
@@
For any distribution $P_{1}(\bm{x})$ in $\epsilon$-neighborhood of $P_{0}(\bm{x})$ with $h(\bm{x})$ satisfying $\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\,\,|h(\bm{x})|\leq 1$, we have
@@
VI-B Realization of Quasi-$\varepsilon$-neighborhood Distributions
@@
In the following, we reconsider the distribution family $P_{n}^{(n)}\sim\mathcal{N}(\bm{0},\mu\Psi(n)\bm{I}_{n})$ in Section III.
@@
Fig. 4: The metric of weak convergence between $\bar{P}_{n}^{(n)}$ and $\delta_{0}^{n}$. On one hand, the distributions $\bar{P}_{l}^{(m)}$ in the $l$-th row converge to the unique Dirac measure $\delta_{0}^{l}$ as $m\rightarrow\infty$ due to the decreasing variance $P(m)$; on the other hand, the distances between $\bar{P}_{n}^{(n)}$ and $\delta_{0}^{n}$ converge to 00 as $n\rightarrow\infty$.
@@
[…]
@@
From the above analysis, when the reference distribution is the distribution of noise, the truncated Gaussian distributions with decreasing variance are proper convolution operators that will lead to output distributions close to the distribution of noise. Hence, they are legitimate choice for the generating distributions of codebooks for AWGN covert channels. Note that the above conclusions holds for any $10<\tau<$ 1, while for covert communication over AWGN channels, the optimal $\tau$ is evaluated as $\frac{1}{2}$ in Theorem 3 and Theorem 5. The following section will explain the optimality of $\tau=\frac{1}{2}$ by investigating the asymptotic behavior of some divergences between Gaussian distributions.
VII Square Root Law and Approximation of the Distribution of Noise
In Section IV and V, we have witnessed that KL divergence between two Gaussian distributions with zero mean and i.i.d coordinates are the key ingredients in the evaluation of both the achievability and converse bounds of the throughput. In this section, the behavior of several divergences between two Gaussian distributions $\mathbb{P}^{(n)}_{1}\sim\mathcal{N}(\bm{0},(1+\theta_{n})\bm{I}_{n})$ and $\mathbb{P}^{(n)}_{0}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$ is investigated in Subsection A, where $\theta_{n}=c\cdot n^{-\tau}$ representing the decreasing power level. This provides more evidence of the connection between Square Root Law and the convergence behaviors of several divergences when $\tau\geq\frac{1}{2}$ and $\tau<\frac{1}{2}$, respectively. Our analysis provides an explanation why we choose $\Psi(n)=O(\frac{1}{n^{\tau}})$ with $\tau=\frac{1}{2}$ in Gaussian random coding in the achievability.
VII-A Asymptotic behaviors of $\mathbb{V}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})$ and $\mathbb{D}(\mathbb{P}^{(n)}_{1}\|\mathbb{P}^{(n)}_{0})$

### v5

@@ III-C Existence of the Codebooks and Achievability Bound
Our result on the size of key bits is slightly different from the case in discrete memoryless channels [22], and is also different from Theorem 6 in [10] where the numbers of key bits are $O(1)$ if the adversary is less noisy ($D_{Q}\leq D_{P}$). In this setting, the number of key bits also depends on the parameter of truncation $\mu$ with finite blocklength over AWGN channels. Even if the channel between Alice and Willie is severer than the main channel, the number of key bits should be of $O(\sqrt{n})$ if the noise level is only a bit larger. This is the necessary cost of power constraint for the sending codewords with finite blocklength over AWGN channels, which will lead to the truncation for Gaussian random coding in the finite blocklength regime.
Remark 17.
The differences between Theorem 6 here and Theorem 5 in Section V of [12] are as follows. Firstly, the main channel $W_{1}$ and the channel at the adversary $W_{2}$ defined in (5) and (6) here are regarded as the same in [12] (with the same level of background noise $\sigma^{2}$). Hence, the defined $L$ there is $1$, irrespectively of the power level $\sigma^{2}$. On the contrary, the first order and second order asymptotics depend on $\sigma^{2}$ in our setting because we allow the noise levels of (5) and (6) to be different. Intuitively, higher level of background noise at the adversary allows the power level of the sending signal to be higher, which leads to larger capacity. Secondly, our analysis is in the finite blocklength regime, where $n$ can only be sufficiently large and $\epsilon$ is usually a constant such as $10^{-3}$. Thus, we can not utilize $n\rightarrow\infty$ or $\epsilon\rightarrow 0$ as the definition of $L$ in (7) in [12].
We can see that both of the sizes of sub-code and the key depend on $\mu$, $\sigma$ and $\delta$. The first order of $\sigma$ and $\delta$. The first order of $\log M^{*}(n,\epsilon,\sigma,\delta)$ increases with $\mu$, $\sigma^{2}$ and $\delta$, which is consistent with the intuition. With fixed $\mu$ and $\delta$, when the noise level at the channel between Alice and Willie is high, then Alice can increase the power of the code and thus the size of the code without violate the covert constraint and maintain the same reliability. The same reasoning applies for fixed $\mu$ and $\sigma$. For the size of key, we can see that with fixed $\delta$, a larger $\sigma$ will lead to a smaller size of key. This is because the transmitter needs less sub-codes to confuse the adversary if the adversary has higher level of background noise.
III-D Converse Bound
In this section, the converse bound under covert constraint (12) is investigated, and the technique is to reconcile the bound in terms of the output distribution of the channel to get an upper bound of average power level of the codebook, which has been utilized in [25][12]. The difference is that the average power considered here should be over all messages and keys for $n$ channel uses. To prove the converse for any code corresponding to any key, we follow the argument in Theorem 7 of [21] and leverage the pigeon-hole principle in our arguments.
First, following the inequality (231) in [21], we have
$\delta\geq\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})\geq n\mathbb{D}(\bar{Q}^{(n)}_{\bm{Z},\bm{C}}\|Q_{\bm{Z},0}),$ (137)
where $\bar{Q}^{(n)}_{\bm{Z},\bm{C}}=\frac{1}{n}\sum_{i=1}^{n}\sum_{s=1}^{K}\sum_{w=1}^{M}\frac{1}{MK}W_{2}(\bm{z}(i)|\bm{x}_{sw}(i))$ is the average output distribution at the adversary per channel use over all the messages and keys and $Q_{\bm{Z},0}$ is the distribution of noise at the adversary. Following the same arguments around (74) in [12], we have the following lemma.
Lemma 13.
For any $(M,K,n,\epsilon,\Psi)_{a}$ code which satisfies the covert constraint, then the corresponding average power $\Psi$ should be less than or equal to the solution $x$ of the equation
$\mathbb{D}(\bar{\mathbb{Q}}_{G}\|Q_{\bm{Z},0}))=\frac{\delta}{n},$ (138)
where $\bar{\mathbb{Q}}^{(n)}_{G}=\mathcal{N}(\bm{0},(\sigma^{2}+\varpi)\bm{I}_{n}).$
Denote the solution of (138) as $\Psi(\delta,\sigma,n)$. From Lemma 13, it is an upper bound of the average power of any $(M,K,n,\epsilon,\sigma,\delta)$ code . Note that the average power considered here is over all messages and keys with $n$ channel uses. To prove the converse bound for any code, we still need the following corollary.
Corollary 3.
There exists a subset $\mathcal{D}$ of codewords $\bm{x}_{sw},s=1,\cdots,K,w=1,\cdots,M$ of size $MK/n$ such that
$\underset{\|\bm{x}_{sw}\|^{2}\leq n\Psi(n)}{\max}\Psi(n)\leq\frac{n}{n-1}\Psi(\delta,\sigma,n)$ (139)
Proof.
If we sort the power of the codewords from maximum to minimum, let $\mathcal{D}$ be the last $\frac{MK}{n}$ codewords. We claim that these codewords have maximal power less than or equal to $\frac{n}{n-1}\Psi(\delta,\sigma,n)$. Otherwise, all the remaining codewords have power at least $\frac{n}{n-1}\Psi(\delta,\sigma,n)$, then the average power of the whole $MK$ codewords is larger than
$\frac{1}{MK}\underset{\bm{x}\in\mathcal{D}^{c}}{\sum}\frac{n}{n-1}\Psi(\delta,\sigma,n)$
$=$ $\frac{1}{MK}\cdot\frac{(n-1)MK}{n}\cdot\frac{n}{n-1}\Psi(\delta,\sigma,n)$
$=$ $\Psi(\delta,\sigma,n),$ (140)
which is a contradiction. ∎
The above corollary show that there exists “part” of the whole codewords such that the maximal power is bounded. By using existing converse bounds under maximal power constraint, we can upper bound the size of that part, and hence the size of each subcode.
Theorem 7.
For AWGN covert channels with channel models in (5) and (6), we have
$\log$ $M^{*}(n,\epsilon,\sigma,\delta)\leq(n\delta\log e)^{1/2}\sigma^{2}$
$+$ $(2)^{1/2}(n\delta)^{1/4}\sigma(\log e)^{3/4}\cdot Q^{-1}(\epsilon)+O(\log n).$ (141)
Proof.
For any $(M,K,n,\epsilon,\sigma,\delta)$ code, we proceed with the arguments in Corollary 13 and Corollary 3. From the pigeon-hole principle, there exists at least one sub-code $\mathcal{C}^{s}$ such that
$|\mathcal{D}\cap\mathcal{C}^{s}|\geq M/n$ (142)
and all the codewords in $\mathcal{D}\cap\mathcal{C}^{s}$ have maximal power less than $\frac{n}{n-1}\Psi(\delta,\sigma,n)$. Thus, the size of $\mathcal{D}\cap\mathcal{C}^{s}$ must satisfy
$\log|\mathcal{D}\cap\mathcal{C}^{s}|$ $\leq\log M^{*}_{m}(n,\epsilon,\frac{n}{n-1}\Psi(\delta,\sigma,n)),$ (143)
Using a converse bound under maximal power constraint with varying power ( the formula (24) in [23]), we get
$\log$ $|\mathcal{D}\cap\mathcal{C}^{s}|\leq n\log(1+\frac{n}{n-1}\Psi(\delta,\sigma,n))$
- $\sqrt{n\mathsf{V}(\frac{n}{n-1}\Psi(\delta,\sigma,n))}\times\Phi^{-1}(1-\epsilon_{1})+O(\log n),$ (144)
where we have used $\Phi^{-1}(1-\epsilon_{1})$ to substitute [^2] Here $Q(x)=\int_{x}^{\infty}\frac{1}{\sqrt{2\pi}}e^{-t^{2}/2dt}$ $Q^{-1}(\epsilon)$. From the fact that for any $\eta>1$
$\frac{1}{2}[x-\ln(1+x)]\geq\frac{x^{2}}{4\eta},$ (145)
we get an upper bound on $\Psi(\delta,n)$ as $\Psi_{u}(\delta,n)=\sqrt{\frac{4\eta\delta\ln 2}{n}}\cdot\sigma^{2}$ and from (142) and (III-D) we further have
$\begin{split}\log&M^{*}(n,\epsilon,\sigma,\delta)\leq\frac{n}{2}\log\left(1+\frac{n}{n-1}\cdot\sqrt{\frac{4\eta\delta\ln 2}{n}}\cdot\sigma^{2}\right)\\ +&\sqrt{\frac{n\log^{2}e}{2}\cdot V_{2}(n)}\Phi^{-1}(\epsilon)+O(\log n),\end{split}$ (146)
where
$V_{2}(n)=\frac{\sigma^{4}\delta\eta\ln 2+\sqrt{\sigma^{4}\delta\eta n\ln 2}}{\sigma^{4}\delta\eta\ln 2+\sqrt{\sigma^{4}\delta\eta n\ln 2}+\frac{n}{4}}.$ (147)
@@ III-D Converse Bound
Note that both the first and second-order asymptotics of the converse bound in (7) differs from the achievability bound (6) in Theorem 6 on the coefficient $\mu$. It shows that the coding scheme in the achievability is optimal. In practice, the maximal throughput can be obtained by choosing proper parameter $\mu$ depending on the blocklength $n$ and the noise level $\sigma^{2}$ at the adversary.
@@
IV A Local Information Geometry Viewpoint of Gaussian Random Coding for Covert Communication
In Section III, truncated Gaussian distribution is adopted as the generating distribution of the AWGN covert codes. From the arguments in Section III-D, the distribution is optimal in the sense that it leads to nearly maximal coding rate when $n$ is large. In this section, we analyze it from a viewpoint of local information geometry, which helps us to understand the underlying principle of choosing generating distribution for random coding in covert communication.
IV-A Quasi-$\varepsilon$-neighborhood Distributions
@@
For any distribution $P_{1}(\bm{x})$ in $\epsilon$-neighborhood of $P_{0}(\bm{x})$ with $h(\bm{x})$ satisfying $\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\,\,|h_{\varepsilon}(\bm{x})|\leq 1$, assume that $\int_{\mathbb{R}^{n}}p_{0}(\bm{x})h^{n}_{\varepsilon}(\bm{x})d\bm{x}<\infty$ for any $n\geq 2$, we have
@@
IV-B Realization of Quasi-$\varepsilon$-neighborhood Distributions
@@
In the following, we reconsider the distribution family $P_{n}^{(n)}\sim\mathcal{N}(\bm{0},\mu\Psi(n)\bm{I}_{n})$ in Section II-E.
@@
Fig. 3: The metric of weak convergence between $\bar{P}_{n}^{(n)}$ and $\delta_{0}^{n}$. On one hand, the distributions $\bar{P}_{l}^{(m)}$ in the $l$-th row converge to the unique Dirac measure $\delta_{0}^{l}$ as $m\rightarrow\infty$ due to the decreasing variance $P(m)$; on the other hand, the distances between $\bar{P}_{n}^{(n)}$ and $\delta_{0}^{n}$ converge to 0 as $n\rightarrow\infty$.
@@
$\begin{split}&\mathbb{\chi}^{2}(\hat{Q}^{(n)}_{1}\|\hat{Q}^{(n)}_{0})\\ =&\varepsilon^{2}\int_{\mathbb{R}^{n}}h^{2}(\bm{y})\hat{f}_{0}^{(n)}(\bm{y})d\bm{y}\\ \leq&\varepsilon^{2}\cdot\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}\,\,|h^{2}(\bm{y})|\int_{\mathbb{R}^{n}}\hat{f}_{0}^{(n)}(\bm{y})d\bm{y}\\ \leq&\varepsilon^{2}.\end{split}$ (157)
@@
From the above analysis, when the reference distribution is the distribution of noise, the truncated Gaussian distributions with decreasing variance are proper convolution operators that will lead to output distributions close to the distribution of noise. Hence, they are legitimate choice for the generating distributions of codebooks for AWGN covert channels. Note that the above conclusions holds for any $0<\tau<1$, while for covert communication over AWGN channels, the optimal $\tau$ is evaluated as $\frac{1}{2}$ in Theorem 6 and Theorem 7. The following section will explain the optimality of $\tau=\frac{1}{2}$ by investigating the asymptotic behavior of some divergences between Gaussian distributions.
V Square Root Law and Approximation of the Distribution of Noise
In Section III and III-D, we have witnessed that KL divergence between two Gaussian distributions with zero mean and i.i.d coordinates are the key ingredients in the evaluation of both the achievability and converse bounds of the throughput. In this section, the behavior of several divergences between two Gaussian distributions $\mathbb{P}^{(n)}_{1}\sim\mathcal{N}(\bm{0},(1+\theta_{n})\bm{I}_{n})$ and $\mathbb{P}^{(n)}_{0}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$ is investigated in Subsection A, where $\theta_{n}=c\cdot n^{-\tau}$ representing the decreasing power level. This provides more evidence of the connection between Square Root Law and the convergence behaviors of several divergences when $\tau\geq\frac{1}{2}$ and $\tau<\frac{1}{2}$, respectively. Our analysis provides an explanation why we choose $\Psi(n)=O(\frac{1}{n^{\tau}})$ with $\tau=\frac{1}{2}$ in Gaussian random coding in the achievability.
V-A Asymptotic behaviors of $\mathbb{V}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})$ and $\mathbb{D}(\mathbb{P}^{(n)}_{1}\|\mathbb{P}^{(n)}_{0})$